# SpikeSTAG: Spatial-Temporal Forecasting via GNN-SNN Collaboration

### v1

@@ Introduction
Analysis of Table 3 demonstrates that the gating fusion mechanism confers a decisive advantage in temporal forecasting. Across the four benchmarks, the architecture that employs gating to combine LSTM with SSA (W4) uniformly surpasses all alternatives, attaining the highest $R^{2}$ of 0.868, 0.874, 0.973 and 0.987 on METR-LA, PEMS-BAY, Solar and Electricity, respectively, while simultaneously delivering the lowest $RSE$ of 0.375, 0.384, 0.246 and 0.207. The benefit is most pronounced on complex dynamic systems such as traffic data, where the gating mechanism yields an $R^{2}$ improvement of 1.0–1.4% over straightforward concatenation (W3). This gain stems from the gate’s capacity to adaptively balance LSTM’s sequential modeling strength with SSA’s spike-based feature extraction, thereby mitigating feature conflicts.
The ablations also highlight the intrinsic merit of each module. The SSA-only variant (W2) outperforms the LSTM-only one (W1) by 3.1–4.3% on three benchmarks, underscoring the superiority of spiking neural representations for raw temporal signals. The concatenated design (W3) emerges as the second-best choice on periodic datasets such as Solar and Electricity, confirming the complementary value of the dual modules. Notably, on PEMS-BAY the performance of W3 drops below that of W2, revealing the risks of unconstrained feature aggregation.
Energy Reduction
SpikeSTAG simultaneously exploits the temporal dynamics inherent to spiking neural networks and the spatial modeling capacity of graph neural networks, while inheriting the energy-efficient characteristics of SNNs. In this section, we map SpikeSTAG iSpikformer and iTransformer onto 45 nm neuromorphic hardware and derive their theoretical energy expenditure(?). All measurements are obtained by executing inference on the Electricity dataset with a prediction horizon of $L=3$.

### v2

@@ 5.4 Ablation Study
Analysis of Table 3 demonstrates that the gating fusion mechanism confers a decisive advantage in temporal forecasting. Across the four benchmarks, the architecture that employs gating to combine LSTM with SSA (W4) uniformly surpasses all alternatives, attaining the highest $R^{2}$ of 0.868, 0.874, 0.973 and 0.987 on METR-LA, PEMS-BAY, Solar and Electricity, respectively, while simultaneously delivering the lowest $RSE$ of 0.375, 0.384, 0.246 and 0.207. The benefit is most pronounced on complex dynamic systems such as traffic data, where the gating mechanism yields an $R^{2}$ improvement of 1.0â1.4% over straightforward concatenation (W3). This gain stems from the gateâs capacity to adaptively balance LSTMâs sequential modeling strength with SSAâs spike-based feature extraction, thereby mitigating feature conflicts.
The ablations also highlight the intrinsic merit of each module. The SSA-only variant (W2) outperforms the LSTM-only one (W1) by 3.1â4.3% on three benchmarks, underscoring the superiority of spiking neural representations for raw temporal signals. The concatenated design (W3) emerges as the second-best choice on periodic datasets such as Solar and Electricity, confirming the complementary value of the dual modules. Notably, on PEMS-BAY the performance of W3 drops below that of W2, revealing the risks of unconstrained feature aggregation.
5.5 Energy Reduction
SpikeSTAG simultaneously exploits the temporal dynamics inherent to spiking neural networks and the spatial modeling capacity of graph neural networks, while inheriting the energy-efficient characteristics of SNNs. In this section, we map SpikeSTAG iSpikformer and iTransformer onto 45 nm neuromorphic hardware and derive their theoretical energy expenditurehorowitz2014energy . All measurements are obtained by executing inference on the Electricity dataset with a prediction horizon of $L=3$.